OpenStudy (sammietaygreen):

Find the distance between points P1(-1, 1) and P2(4, -4). Then find the coordinates of the midpoint of the segment that has endpoints at P1 and P2.

OpenStudy (caozeyuan):

ok so d=sqrt(a^2+b^2), a=x2-x1,b=y2-y1

OpenStudy (caozeyuan):

here we see x1=-1, x2=4,y1=1,y2=-4

OpenStudy (sammietaygreen):

sqrt 50

OpenStudy (sammietaygreen):

7.1

OpenStudy (caozeyuan):

midpoint has formula((x1+x2)/2,(y1+y2)/2)

OpenStudy (sammietaygreen):

Do you input those numbers into that for the second answer?

OpenStudy (caozeyuan):

yes

OpenStudy (sammietaygreen):

would it be (3/2, -3/2)? but it looks like they're asking for it in decimal form

OpenStudy (caozeyuan):

well, 3/2 is 1.5 right?

OpenStudy (sammietaygreen):

yes so it would be 1.5, -1.5

OpenStudy (sammietaygreen):

so the entire answer would consist of 7.07; (1.5, -1.5)

OpenStudy (caozeyuan):

great! you did it!

OpenStudy (sammietaygreen):

thank you!

OpenStudy (caozeyuan):

np